OpenStudy (anonymous):

Suppose AD=MxM identity Matrix. Show that any vector b in space R Mthe equation Ax=b where b and x are vectors.

OpenStudy (anonymous):

i can not wait for this response. i am sitting on the edge of my seat

OpenStudy (turingtest):

i'm not seeing a question here

OpenStudy (turingtest):

show that any vector in \(\mathbb R^m\) satisfies \(A\vec x=\vec b\) if \[AB_{m\times m}=I\]?

OpenStudy (anonymous):

its a proof. Show that any vector b in \[\mathbb{R}^m\] the equation ax=b

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

or you have \[AD=I_{m\times m}\]correct?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

and there is a hint saying think about the eq ADb=b

OpenStudy (turingtest):

oh ok that may help

OpenStudy (anonymous):

show that any vector in Rm satisfies Ax⃗ =b⃗ if ABm×m=I this is correct

OpenStudy (turingtest):

ok i understand the problem i think but i'm rusty lol

OpenStudy (anonymous):

This is what i had ADb=b ADAx=b IAx=b Ax=b but i used the fact that Ax=b which is what i needed to prove in step 2

OpenStudy (anonymous):

we also know that AD=Ad1+Ad2....Adn where di is the columns of D. from the definition of matrix multiplication

OpenStudy (anonymous):

and if that equals the identity does that imply that Ad1=e1 and so forth?

OpenStudy (anonymous):

where e1= equals the idenity matrix of column 1

OpenStudy (turingtest):

i suppose it would yeah

OpenStudy (turingtest):

what about something like\[AB\vec b=\vec b\\\vec b=B^{-1}A^{-1}\vec b=\vec x\\A\vec b=\vec x\]i don't know if that is less tautological than yours though

OpenStudy (anonymous):

I think i got it. To show that all Ax=b for all b we need to show that Ax is an onto function. To show some thing is onto we need to show that the columns of A span Rm which is done by showing AD=I

OpenStudy (anonymous):

o okay yours makes sense as well

OpenStudy (turingtest):

hm you are going beyond what i remember from LA well lol, but i think it works if you just assert that \(\vec b\) can be \(\vec x\) since the inverse of the identity is the identity

OpenStudy (turingtest):

in the first part of the proof i mean

OpenStudy (turingtest):

maybe take out the x though, hm still seems tautological

OpenStudy (anonymous):

alright sounds good

OpenStudy (turingtest):

no it's all wrong lol i need to rethink it

OpenStudy (anonymous):

haha alright

OpenStudy (turingtest):

yeah way too rusty on my LA.... dang maybe @AccessDenied or @thomaster can help?

OpenStudy (turingtest):

remove my medal :P

OpenStudy (anonymous):

ill leave it thanks for the help

OpenStudy (turingtest):

hehe thanks for the reminder i need to review this stuff

OpenStudy (turingtest):

oh dude i think i got it

OpenStudy (anonymous):

alright sounds good haha

OpenStudy (turingtest):

dang it false alarm lol

OpenStudy (anonymous):

its fine i am sure that it will be okay.

OpenStudy (turingtest):

i think that it is enough to say that\[AB=I_{m\times m}\implies B=A^{-1}\]so you have\[A^{-1}AA^{-1}\vec b=A^{-1}\vec b=\vec x\]which must be in \(\mathbb R^m\) because \(A\) is an \(m\times n\) matrix, and since \(A\) is invertible, for every \(\vec x\) in \(\mathbb R^m\) there must be a unique \(\vec b\) such that\[A\vec x=\vec b\]

OpenStudy (anonymous):

yep that works. because you showed that A is onto and 1 to 1. Thanks!

OpenStudy (turingtest):

yw :)